# Synthesizing Moving People with 3D Control

### v1

@@ Abstract
Figure 1: The Imitation Game: Given a video of a person "The Actor", we want to transfer their motion to a new person "The Imitator". In this figure, the first row shows a sequence of frames of the actor (Michelle Kwan), doing her Olympics ’98 performance. The inset row shows the 3D poses extracted from this video. Now, given any single image of a new person The Imitator, our model can synthesize new renderings of the imitator, to copy the actions of the actor in 3D.
@@ 1 Introduction
Given a random photo of a person, can we accurately animate that person to imitate someone else’s action? This problem requires a deep understanding of how human poses change over time, learning priors about human appearance and clothing. For example, in Figure 1 the Actor can do a diverse set of actions, from simple actions such as walking and running to more complex actions such as fighting and dancing. For the Imitator, learning a visual prior about their appearance and clothing is essential to animate them at different poses and viewpoints. To tackle this problem, we propose 3DHM, a two-stage framework (see Figure 2) that synthesizes 3D Human Motions by completing a texture map from a single image and then rendering the 3D humans to imitate the actions of the actor.
@@ 2 Related Works
Controllable Human Generation. Human generation is not an easy task. Unlike image translation [16], generating different humans requires the model to understand the 3D structure of the human body. Given arbitrary text prompts or pose conditions [5, 15], we often find out that existing generative models often generate unreasonable human images or videos. Diffusion-HPC [31] proposes a diffusion model with Human Pose Correction and finds that injecting human body structure priors within the generation process could improve the quality of generated images. ControlNet [34] is designed on neural network architecture to control pre-trained large diffusion models to support additional input conditions, such as Openpose [6]. GestureDiffuCLIP [3] designs a neural network to generate co-speech gestures. However, these techniques are not tailored for animating humans, which cannot guarantee the required human appearance and clothing.
Synthesizing Moving People. Synthesizing moving people is very challenging. For example, Make-a-Video [24] or Imagen Video [23] could synthesize videos based on a given instruction. However, the generated video cannot accurately capture human properties correctly and may cause the weird composition of generated humans. Prior methods [8, 29] learn pose-to-pixels mapping directly. However, these designs could only be trained and used for one person. Recent works such as SMPLitex [7] consider human texture estimation from a single image to animate a person. However, there is a visual gap between rendered people via predicted texture map and real humans. Many works start to directly predict pixels based on diffusion models, such as Dreampose [14] and DisCO [28]. DreamPose is controlled by DensePose [10], it aims to synthesize a video containing both human and fabric motion based on a sequence of human body poses. DisCO is directly controlled by Openpose [6], and it aims to animate the human based on the 2D pose information. However, the approach of aligning output pixels for training regularization often leads these models to become overly specialized to certain training data. Moreover, this methodology limits the models’ generalization capabilities, as they often perform well on a few people whose data distribution closely matches that of the training dataset.
Figure 2: Overview of 3DHM: we show an overview of our model pipeline. Given an image of the imitator and a sequence of 3D poses from the actor, we first generate a complete full texture map of the imitator, which can be applied to the 3D pose sequences extracted from the actor to generate texture-mapped intermediate renderings of the imitator. Then we pass these intermediate renderings to the Stage-2 model to project the SMPL mesh rendering to more realistic renderings of real images. Note: red boxes represent inputs, yellow boxes represent intermediate predictions from stage 1, and blue boxes represent the final outputs from stage 2. To create a moving person animation with variable duration and any number of 3D poses, it is only necessary to execute stage 1 once in order to acquire a complete texture map.
@@ 3 Synthesizing Moving People
Our approach 3DHM is composed of two core parts: Inpainting Diffusion for texture map in-painting as Stage-1 and Rendering Diffusion for human rendering as Stage-2. Figure 2 shows a high-level overview of our framework. In Stage-1, first, for a given single view image, we extract a rough estimate of the texture map by rendering the meshes onto the image and assigning pixels to each visible mesh triangle such that when rendered again it will produce a similar image as the input image. This predicted texture map has only visible parts of the input image. The Stage-1 Diffusion in-painting model takes this partial texture map and generates a complete texture map including the unseen regions. Given this completed texture map, we generate intermediate renderings of SMPL [17] meshes and use Stage-2 model to project the body-tight renderings to more realistic images with clothing. For the Stage-2 Diffusion model, we apply 3D control to animate the imitator to copy the actions of the actor.
@@ 3.1 Texture map Inpainting
Target. Since the objective of this modeling is to generate complete texture maps, we generate a pseudo-complete texture map using video data. Since the 4DHumans can track people over time, it continually updates its internal texture map representations as a moving average of visible regions. However to produce more sharp images, for the generative task we found that a median filtering is more suitable than a moving average. While this technique can be applied to any video, in this stage we rely on 2,205 human videos. For each human video, we first extract a partial texture map from each frame. Since each video contains 360 degrees of human views, we calculate a pseudo-complete texture map from a whole video and set it as the target output for Stage 1. In detail, we take the median overall visible parts of texture maps of a video.
Figure 3: Stage-1 of 3DHM: In the first stage, given a single view image of an imitator, we first apply 4Dhumans [9] style sampling approach to extract partial texture map and its corresponding visibility map. These two inputs are passed to the in-painting diffusion model to generate a plausible complete texture map. In this example, while we only see the back view of the imitator, the model was able to hallucinate a plausible front region that is consistent with their clothing.
Model. We finetune directly on the Stable Diffusion Inpainting model [21] that shows great performance on image completion tasks. We input a partial texture map and corresponding visibility mask and obtain the recovered predicted map for the human. We lock the text encoder branch and always feed ‘real human’ as input text of fixed Stable Diffusion models. We refer to our trained model as Inpainting Diffusion. See Figure 3 for the model architecture.
@@ 3.2 Human Rendering
In Stage 2, we aim to obtain a realistic rendering of a human imitator doing the actions of the actor. While the intermediate renderings (rendered with the poses from the actor and texture map from Stage-1) can reflect diverse human motion, these SMPL mesh renderings are body-tight and cannot represent realistic rendering with clothing, hairstyles, and body shapes. For example, if we input a scene where a girl is wearing a dress and she is dancing, the intermediate renderings might be able to “dance" but it is impossible to animate the skirt with SMPL mesh rendering. To train this model, in a fully self-supervised fashion, we assume the actor is the imitator, after all a good actor should be a good imitator. This way, we can take any video, and get a sequence of poses from 4DHumans [9] and take any single frame, and get a complete texture map from Stage-1, then get the intermediate renderings by rendering the texture maps on the 3D poses. Now, we have paired data of intermediate renderings and real RGB images. Using this, we collect a large amount of paired data and train our Stage-2 diffusion model with conditioning.
Input: We first apply the generated texture map (fully complete) from Stage 1 to actor 3D body mesh sequences to an intermediate rendering of the imitator performing the actions of the actor. Note at this time, intermediate rendering can only reflect the clothing that fits the 3D mesh (body-tight clothing) but fails to reflect the texture outside the SMPL body, such as the puffed-up region of a skirt, winter jacket, or hat. To obtain the human with complete clothing texture, we input the obtained intermediate renderings and the original image of the person into Rendering Diffusion to render the human in a novel pose with a realistic appearance.
Figure 4: Stage-2 of 3DHM: This figure shows the inference of our Stage-1 approach. Given an intermediate rendering of the imitator with the pose of the actor and the actual RGB image of the imitator, our model can synthesize realistic renderings of the imitator on the pose of the actor.
@@
Model. Similar to ControlNet, we directly clone the weights of the encoder of the Stable Diffusion [20] model as our Controllable branch ("trainable copy") to process 3D conditions. We freeze the pre-trained Stable Diffusion and input noisy latents ($64\times 64$). In the meanwhile, we input a texture mapped 3D human at time $t$ and original human photo input into a fixed VAE encoder and obtain texture mapped 3D human latents ($64\times 64$) and appearance latents ($64\times 64$) as conditioning latents. We feed these two conditioning latents into Rendering Diffusion Controllable branch. The key design principle of this branch is to learn textures from human input and apply them to the texture mapped 3D human during training through the denoising process. The goal is to render a real human with vivid textures from the generated(texture mapped) 3D human from Stage 1. We obtain the output latent and process it to the pixel space via diffusion step procedure and fixed VAE decoder. Same to Stage 1, we lock the text encoder branch and always feed ‘a real human is acting’ as input text of fixed Stable Diffusion models. We refer to our trained model as Rendering Diffusion. In Rendering Diffusion, we predict outputs frame by frame. We show the Stage 2 workflow in Figure 4.
@@ 4.1 Experimental Setup
Evaluation Metrics. We evaluate the quality of generated frames of our method with image-based and video-based metrics. For image-based evaluation, we follow the evaluation protocol of DisCO [28] to evaluate the generation quality. We report the average PSNR [13], SSIM [30], FID [12], LPIPS [35], and L1. For video-based evaluation, we use FVD [26]. For pose evaluating 3D pose accuracy we use MPVPE and PA-MVPVE. MPVPE [18], or Mean Per-Vertex Position Error, is a critical metric in 3D human pose estimation, which quantifies the average distance between predicted and actual 3D vertices across a model. This measurement is essential for evaluating the accuracy of 3D reconstructions and pose estimations, with a lower MPVPE indicating higher precision. Complementing this, PA-MPVPE, or Procrustes-Aligned Mean Per-Vertex Position Error, adds another dimension to this evaluation. It involves aligning the predicted and ground truth data using Procrustes Analysis, which neutralizes differences in orientation, scale, and position before calculating the mean error. This alignment allows PA-MPVPE to focus on the structural accuracy of predictions, making it a valuable metric for assessing the relative positioning of vertices in a model, independent of their absolute spatial coordinates.
Implementation Details. As for training all the datasets, we set the constant learning rate as 5e-05 and use the pre-trained diffusion models from diffusers [27] for both Stage-1 and Stage-2. As for Stage 1 Inpainting Diffusion, we finetune on Stable Diffusion Inpainting models [21], which has an 859M total number of trainable parameters and 206M total number of non-trainable parameters, since the VAE is frozen during this stage. We train Rendering Diffusion for 50 epochs and it takes about 2 weeks to run our model on our soup of training datasets. As for Stage 2 Rendering Diffusion, we train the Controllable branch and freeze Stable Diffusion backbones. The total number of trainable parameters in this case is 876M and the total number of non-trainable parameters is 1.1B. We train Rendering Diffusion for 30 epochs and it takes about 2 weeks to run our model on training datasets based on 8 NVIDIA A100 GPUs with a batch size of 4. As for inference, we only need to run Stage-1 once to reconstruct the full texture map of the imitator, and it is used for all other novel poses and viewpoints. We run Stage-2 inference for each frame independently, however since the initial RGB frame of the imitator is conditioned for all frames, the Stage-2 model is able to produce samples that are temporarily consistent.

### v2

@@ Abstract
Figure 1: The Imitation Game: Given a video of a person "The Actor", we want to transfer their motion to a new person "The Imitator". In this figure, the first row shows a sequence of frames of the actor from a ballerina Dance of the Sugar Plum Fairy. The inset row shows the 3D poses extracted from this video. Now, given any single image of a new person The Imitator, our model can synthesize new renderings of the imitator, to copy the actions of the actor in 3D.
@@ 1 Introduction
Given a random photo of a person, can we accurately animate that person to imitate someone else’s action? This problem requires a deep understanding of how human poses change over time, learning priors about human appearance and clothing. For example, in Figure 1 the Actor can do a diverse set of actions, from simple actions such as walking and running to more complex actions such as fighting and dancing. For the Imitator, learning a visual prior about their appearance and clothing is essential to animate them at different poses and viewpoints. To tackle this problem, we propose 3DHM, a diffusion framework (see Figure 2) that synthesizes 3D Human Motions by completing a texture map from a single image and then rendering the 3D humans to imitate the actions of the actor.
Figure 2: Overview of 3DHM: we show an overview of our model pipeline. Given an image of the imitator and a sequence of 3D poses from the actor, we first generate a complete full texture map of the imitator, which can be applied to the 3D pose sequences extracted from the actor to generate texture-mapped intermediate renderings of the imitator. Then we pass these intermediate renderings to the Stage-2 model to project the SMPL mesh rendering to more realistic renderings of real images.
@@ 2 Related Works
Controllable Human Generation. Human generation is not an easy task. Unlike image translation [18], generating different humans requires the model to understand the 3D structure of the human body. Given arbitrary text prompts or pose conditions [4, 17], we often find out that existing generative models often generate unreasonable human images or videos. Diffusion-HPC [36] proposes a diffusion model with Human Pose Correction and finds that injecting human body structure priors within the generation process could improve the quality of generated images. ControlNet [41] is designed on neural network architecture to control pre-trained large diffusion models to support additional input conditions, such as Openpose [5]. GestureDiffuCLIP [2] designs a neural network to generate co-speech gestures. However, these techniques are not tailored for animating humans, and do not guarantee the required human appearance and clothing.
Synthesizing Moving People. Synthesizing moving people is very challenging. For example, Make-a-Video [29] or Imagen Video [28] could synthesize videos based on a given instruction. However, the generated video cannot accurately capture human properties correctly and may cause the weird composition of generated humans. Older methods [7, 33] learn pose-to-pixels mapping directly, but they require being trained for each new person separately. Recent works such as SMPLitex [6] consider human texture estimation from a single image to animate a person. However, there is a visual gap between rendered people via predicted texture map and real humans. Many works start to directly predict pixels based on diffusion models, such as Dreampose [16], DisCO [32], AnimateAnyone [15], MagicAnimate [38], Champ [43], etc. [8, 34, 21, 23]. DreamPose and MagicAnimate is controlled by DensePose [10], it aims to synthesize a video containing both human and fabric motion based on a sequence of human body UV or Segmentation maps. DisCO and AnimateAnyone is directly controlled by Openpose [5], and it aims to animate the human based on the 2D pose information. Champ [43] utilizes the multiple condition maps rendered from SMPL mesh to further enhance detailed controllability. However, the approach of aligning output pixels for training regularization often leads these models to become overly specialized to certain training data. Moreover, this methodology limits the models’ generalization, as they often perform well on a few people whose data distribution closely matches that of the training dataset.
@@ 3 Synthesizing Moving People
Our approach 3DHM is composed of two core parts: Inpainting Diffusion for texture map in-painting as Stage-1 and Rendering Diffusion for human rendering as Stage-2. Figure 2 shows a high-level overview of our framework. In Stage-1, first, for a given single view image, we extract a rough estimate of the texture map by rendering the meshes onto the image and assigning pixels to each visible mesh triangle such that when rendered again it will produce a similar image as the input image. This predicted texture map has only visible parts of the input image. The Stage-1 Diffusion in-painting model takes this partial texture map and generates a complete texture map including the unseen regions. Given this complete texture map, we generate intermediate renderings of SMPL [20] meshes and use Stage-2 model to project the body-tight renderings to more realistic images with clothing. For the Stage-2 model, we apply 3D control to animate the imitator to copy the actions of the actor.
@@ 3.1 Texture map Inpainting
Target. We train our model on a large 3D human texture dataset [19], which contains 50k high-fidelity textured UV map of SMPL [20]. To strengthen the model’s 3D geometry consistency in completing the partial texturemap, We densely sample a group of visibility masks from 360 degrees of SMPL mesh, which then mask out Ground-Truth texture map to produce the pseudo-partial texture map during training the inpainting model. Benefiting from the extensive collection of texture maps from diverse human appearances, as well as the numerous visibility masks from various viewpoints.
Figure 3: Stage-1 of 3DHM: In the first stage, given a single view image of an imitator, we first apply 4Dhumans [9] style sampling approach to extract partial texture map and its corresponding visibility map. These two inputs are passed to the in-painting diffusion model to generate a plausible complete texture map. In this example, while we only see the front view of the imitator, the model was able to hallucinate a plausible back region that is consistent with their clothing.
Model. We finetune directly on the Stable Diffusion Inpainting model [26] that shows great performance on image completion tasks. Given a single RGB human image, we predict the human mesh and calculate its corresponding visibility mask and partial texture map, which is then recovered by the in-painting model to complete texture map for the human. We lock the text encoder branch during training and feed "3D realistic human, UV texturemap" as input text condition. We refer to our trained model as Inpainting Diffusion. See Figure 3 for the model architecture.
@@ 3.2 Human Rendering
In Stage 2, we aim to obtain a realistic rendering of a human imitator doing the actions of the actor. While the intermediate renderings (rendered with the poses from the actor and texture map from Stage-1) can reflect diverse human motion, these SMPL mesh renderings are body-tight and cannot represent realistic rendering with clothing, hairstyles, and body shapes. We train a model for realistic rendering, in a fully self-supervised fashion, by relying on the actor as the imitator. We obtain a sequence of poses from 4DHumans [9] for each training video and use Stage-1 on single frames to obtain a complete texture map. We then pair the intermediate renderings (i.e. the rendered texture maps on the 3D poses) with the original frames from which they were obtained. We collect a large amount of paired data and train our Stage-2 diffusion model with conditioning.
Input: We first apply the generated complete texture map from Stage-1 to the actor’s 3D body mesh sequences to obtain the intermediate rendering. Note that the rendering can only reflect the clothing that fits the 3D mesh (body-tight clothing) but fails to reflect the texture outside the SMPL body (e.g., the puffed-up skirt region, or hat). To obtain the human with complete clothing texture, we input the obtained intermediate renderings and the original image of the person into Rendering Diffusion to render the human in a novel pose with a realistic appearance.
@@
Model. Similar to ControlNet, we directly clone the weights of the encoder of the Stable Diffusion [25] model as our Controllable branch ("trainable copy") to process 3D conditions. We freeze the pre-trained Stable Diffusion. In the meanwhile, we input a texture-mapped 3D human at time $t$ and original human photo input into a fixed VAE encoder and obtain texture-mapped 3D human latents ($64\times 64$) and appearance latents ($64\times 64$) as conditioning latents. We feed these two conditioning latents into Rendering Diffusion Controllable branch. The key design principle of this branch is to learn textures from human input and apply them to the texture-mapped 3D human during training through the denoising process. The goal is to render a real human with vivid textures from the generated(texture-mapped) 3D human from Stage 1. We obtain the output latent and process it to the pixel space via diffusion step procedure and fixed VAE decoder. We refer to our trained model as Rendering Diffusion. In Rendering Diffusion, we predict outputs frame by frame. We show the Stage 2 workflow in Figure 4.
Figure 4: Stage-2 of 3DHM: Given an intermediate rendering of the imitator with the pose of the actor and the actual RGB image of the imitator, our model can synthesize realistic renderings of the imitator on the pose of the actor.
4 A Complete Approach
In this section, we discuss how to scale up our method to real-world domains. We first discuss the challenges of visual appearance injection for Diffusion models and then propose a novel framework to enhance consistent appearance alignment. For any given reference human images (the imitators), our network can generate high-fidelity results benefiting from the enhanced appearance encoder [39]. To ensure the visual consistency of both human identity and background from reference images, we use a trainable copy of base Stable Diffusion model [27] to inject appearance information that perfectly aligns with the backbone denoising Diffusion model. To encourage temporally smooth reconstructions, we utilize a temporal diffusion model [11] to learn the temporal correlation within motion sequences. To generate smooth and consistent videos, we propose an easy but efficient method that takes the previous frame for consequential video generation. The detailed framework is shown in Figure 5.
4.1 Enhance Appearance Alignment
The key challenge in scaling up our method to real-world domains is to both maintain the complex background and the human appearance from reference images consistently. The Stable Diffusion Model [27] is trained on text-to-image tasks, and focuses on semantics of the generated image. However, in our Stage-2 rendering task, instead of semantic features, the model needs more low-level visual features to reconstruct the detailed structure and appearance from the input images. Given the imitator’s images as reference, our approach simultaneously leverages the capabilities of Stable Diffusion and uses the reference image prompts to obtain more accurate generation. We use a lightweight image adapter [39] to condition diffusion on image prompts. We also add a trainable branch of the Stable Diffusion model, namely Reference Net, to enhance appearance consistency on both the input image’s background and human appearance.
Input. Same with Stage-2, we input the generated imitator’s texture map from Stage-1 with actor’s 3D mesh sequences to get the intermediate rendering. Then the intermediate rendering is input to the 3D Controllable branch as motion condition. The original imitator’s RGB image is input to the Reference Net and the image adaptor as the appearance guidance.
Target. We scale up our model on about 1000 real human videos collected from the Internet, each containing 2-10 seconds solo dancing. For this stage, we keep the assumption that: imitator and actor are identical and randomly sample the imitator and actor. The objective now is to drive the reference imitator image with actor’s pose to generate the corresponding target image. We trained on our real and virtual datasets together to teach model focus on complex pose and appearance variance, and the 3D view consistency respectively.
Figure 5: Scaled up Stage-2 of 3DHM Model: To enable consistent background and human generation, we train ReferenceNet with ControlNet, and then only finetune the temporal-attention layer of the UNet and keep other components frozen.
Model. To better align the input image’s appearance with the denoising backbone, we make a trainable copy of the pre-trained Stable Diffusion as our ReferenceNet. Notice that different from the Section 3.2, instead of adding appearance latents with 3D human latents together, we now separately input the imitator’s appearance latents to ReferenceNet to extract level of details appearance latents. The appearance latents are then injected into the denoising backbone to condition the denoising process, which help to keep consistent background and human appearance for different poses. The pretrained IP-adapter [39] is also integrated into all branches to keep human identity.
4.2 Temporal Consistency
Once our image-to-image model learned to generate the imitator frames we can apply it frame-by-frame for image-to-video generation. However, the generated video may suffer from jittering due to the lack of temporal consistency. Based on the Stage-2 model, we adapt a temporal model pretrained on a large video dataset  [11] to learn the motion’s coherence and appearance consistency. Though previous works  [32, 16, 15] also have similar temporal layers and success for short video, they still suffer from inconsistency in long video generation results. To effectively solve the unstable and randomness between each short video clip generated from temporal sliding windows, we design an easy process to concatenate the previous frame’s latents with the consequential video.
Input. In this stage, our model is optimized directly on video data. During training, for each video clip, we extract $F$ consecutive frames as the target of actor’s motion sequence and randomly pick a frame as the imitator’s reference image. Now the 3D Control branch takes $F$ consecutive intermediate rendering as motion sequences to drive the imitator’s image to generate a temporal coherent video. Notice that our Reference Net here will not cost extra computing time since there is only one reference image for each video clip.
Model. The short video clip is now defined as $V\in\mathbb{R}^{B\times C\times F\times H\times W}$, with batch size $B$, the number of channels $C$, the number of consecutive frames $F$, height $H$ and width $W$ respectively. The temporal layers are inserted at each resolution level. For level $i$ , the 5D latents $v_{i}\in\mathbb{R}^{b\times c\times f\times h\times w}$ is reshaped to ${(b\times h\times w)\times f\times c}$ within the temporal layer for self-attention to align feature maps across frames. The temporal attention mechanism above not only effectively smooths the flickering and jittering, but also improves the motion and appearance consistency in generated videos. However, during long video generation, since the video clips are generated independently and concatenated together, the different random noises will cause inconsistency between each video clip. To facilitate the cross-clips consistency, we take the last frame $V_{k}^{f}$ from $k$-th generated video clip $V_{k}^{1:f}$ to condition on the next video clip generation $V_{k+1}^{1:f}$. The $V_{k}^{f}$ is input to the Reference Net to extract corresponding latents for each resolution level $i$, and then fed into the temporal layers concatenated with original 5D latents along the temporal dimension. The conditioned temporal layers at level $i$ now attention across latent $v_{i}^{0:f}\in\mathbb{R}^{b\times c\times(1+f)\times h\times w}$ and then trunck the previous frame $v_{i}^{0}$ to get the conditioned results $\bar{v}_{i}^{1:f}$. During this stage, we freeze all other parameters and train only the temporal model. Zero-initialize [41] is also applied to the temporal layers to eliminate harmful noise during training.
@@ 5.1 Experimental Setup
Evaluation Metrics. We evaluate the quality of generated frames of our method with image-based and video-based metrics. For image-based evaluation, we follow the evaluation protocol of DisCO [32] to evaluate the generation quality. We report the average PSNR [14], SSIM [35], FID [13], LPIPS [42], and L1. For video-based evaluation, we use FVD [30]. For pose evaluating 3D pose accuracy we use Mean Per-Vertex Position Error (MPVPE) and Procrustes-Aligned Mean Per-Vertex Position Error (PA-MVPVE [22]).
Implementation Details. We set a learning rate of 5e-05 and use the pre-trained diffusion models for both stages. For Stage 1 Inpainting Diffusion, we fine-tune Stable Diffusion Inpainting models [26] We train Rendering Diffusion for 50 epochs (requires about 2 weeks on our compute). For Stage 2 Rendering Diffusion, we train the Controllable branch and freeze Stable Diffusion backbones. The total number of trainable parameters in this case is 876M. We train Rendering Diffusion for 30 epochs (requires about 2 weeks on 8 NVIDIA A100 GPUs with a batch size of 4). During inference, we only need to run Stage-1 once to reconstruct the full texture map of the imitator, which is used for all other novel poses and viewpoints. We run Stage-2 inference for each frame independently, however since the initial RGB frame of the imitator is conditioned for all frames, the Stage-2 model is able to produce samples that are temporarily consistent.
(a) 3DHM with random 3D poses from various viewpoints. We show that even if the person’s photo is from a side angle, our stage 1 can help reconstruct the full texture map, which could be used to obtain full body information. Stage 2 can add texture information based on a given input.
(b) 3DHM with motions from random YouTube Videos.
Figure 6: Qualitative results on different viewpoints of the same pose and motions from random videos.